OpenStudy (anonymous):

You have already invested $600 in a stock with an annual return of 11%. How much of an additional $1,400 should be invested at 12% and how much at 6% so that the total return on the entire $2,000 is 9%? (Round each answer to the nearest cent.)

hero (hero):

Your initial equation where x = 600: 600 + y + z = 2000 600(.11) + .12y + .6z = 2000(.09) Now, after reducing both equations you get: y + z = 2000 - 600 .12y + .06z = (2000)(.09) - 600(.11) Leaving you with a simpler equation of y + z = 1400 .12y + .06z = 114 Now, finish solving the system. Good luck.